# Isotropic Covariance Functions for Disparate Spatial Resolutions

### v2

@@ Abstract
Distances between sets arise naturally in data fusion problems involving both point-referenced and areal observations, as well as in set-indexed stochastic processes more broadly. However, commonly used constructions of distances on sets, including those derived from the Hausdorff distance, generally fail to be conditionally negative definite, precluding their use in isotropic covariance models. We propose the ball–Hausdorff distance, defined as the Hausdorff distance between the minimum enclosing balls of bounded sets in a metric space. For length spaces, we derive an explicit representation of this distance in terms of the associated centers and radii. We show that the ball–Hausdorff distance is conditionally negative definite whenever the underlying metric is conditionally negative definite, which implies, via Schoenberg’s theorem, an isometric embedding into a Hilbert space. As a consequence, broad classes of isotropic covariance functions, including the Matérn and powered exponential families, are valid for random fields indexed by sets. The resulting construction reduces set-to-set dependence to low-dimensional geometric summaries, leading to substantial computational simplifications in covariance evaluation.
Keywords: Change of support, Geostatistics, Hausdorff distance, Length spaces, Random fields, Spatial data fusion
@@ 1 Introduction
Many modern spatial analyses involve observations recorded over spatial units of differing locations, sizes, and geometries, creating fundamental challenges for modeling dependence across spatial supports. Such settings are common in environmental science and related areas such as air pollution epidemiology (e.g., Ellis-Soto et al.,, 2023; Zhong et al.,, 2024), where different aspects of the same underlying process are observed at multiple spatial resolutions. Several closely related statistical problems arise from this mismatch. Change of support refers to predicting a process indexed on one spatial scale using observations collected on another (Gelfand et al.,, 2001), with downscaling denoting the special case in which information aggregated over coarse spatial units is used to infer values at a finer resolution (Zheng et al.,, 2025). When observations on the same phenomenon are simultaneously available at multiple spatial resolutions, the problem is commonly framed as spatial data fusion (Moraga et al.,, 2017). In addition, response and explanatory variables are often observed on incompatible spatial supports, a setting known as spatial misalignment (Gotway and Young,, 2002). Despite differing terminology, these problems share a common structure: the need to model dependence between spatial units that differ in their spatial support.
@@
Most existing approaches to multi-resolution spatial data model dependence by treating observations at different supports as aggregations of a common underlying continuous process, which leads to substantial computational and theoretical limitations. Areal measurements are typically represented as integrals of an underlying stochastic process (Fuentes and Raftery,, 2005). However, the implied finite-dimensional probability density functions involve stochastic integrals that are analytically intractable even when the underlying process is a Gaussian Process (GP). Consequently, inferences rely on numerical approximations (Cowles et al.,, 2002) or Monte Carlo integration (Gelfand et al.,, 2001), which can be computationally demanding and introduce hard-to-quantify bias (Gonçalves and Gamerman,, 2018). An alternative strategy avoids aggregation by defining stochastic processes directly on collections of spatial sets. The Hausdorff distance (Hausdorff,, 2005, pg. 167) has been explored as a basis for such constructions (Godoy et al.,, 2025), but its direct construction does not guarantee positive definite (PD) functions, leaving the existence and validity of the induced process unresolved. This lack of theoretical guarantee, coupled with high computational costs (Knauer et al.,, 2011), necessitates a more robust framework for set-to-set dependence.
@@
Constructing valid PD functions becomes substantially more delicate once the underlying index space departs from Euclidean geometry. While many PD functions are known for $\mathbb{R}^{p}$ equipped with the Euclidean distance (Cressie,, 1993, Ch. 2.5.1), their positive definiteness may not extend to other spaces (Li et al.,, 2023). In fact, a function that is PD on $\mathbb{R}^{1}$ may not remain PD on $\mathbb{R}^{2}$, although the reverse is true (Cressie,, 1993, p. 22). This geometric sensitivity has motivated recent work on characterizing PD covariance functions on non-Euclidean spaces. The particular case of the sphere has been extensively studied (Banerjee,, 2005; Gneiting,, 2013; Guinness and Fuentes,, 2016; Porcu et al.,, 2021). Certain covariance functions exhibit the notable property of being PD on both the (p1)(p-1)-dimensional sphere ($\mathbb{S}^{p-1}$) and $p$-dimensional Euclidean space (Porcu et al.,, 2016). More recently, analogous questions have arisen for covariance functions indexed by graphs and their edges (Anderes et al.,, 2020). Despite these advances, there is no general construction of distances on spatial sets that both ensures positive definiteness of isotropic covariance functions and avoids the computational burdens associated with aggregation-based models.
@@
To address this gap, we propose a framework for constructing isotropic covariance functions directly on collections of spatial sets, which accommodates point-referenced, areal, and mixed spatial data within a unified formulation. The goal is to circumvent the analytical and computational intractability of aggregation-based models while ensuring mathematical validity. To this end, we define a pseudometric, the ball-Hausdorff distance, based on the Hausdorff distance between minimum enclosing balls of sets. Under mild assumptions, we derive a closed-form expression for the proposed distance and show that it is conditionally negative definite (CND) whenever the underlying domain metric is CND. By Schoenberg’s theorem (Schoenberg,, 1938), this implies that the square-root of the ball-Hausdorff distance admits an isometric embedding into a Hilbert space, thereby providing a rich class of valid isotropic covariance functions indexed by spatial sets. As we demonstrate in numerical studies, this theoretical validity is a critical distinction, as the standard Hausdorff distance frequently yields invalid covariance structures in identical settings.
@@ 2 Preliminaries
Let $D$ denote an index set and $d\,:\,D\times D\to[0,\infty)$ a distance function. The pair $(D,d)$ is a pseudometric space if $d$ satisfies all the metric axioms except the identity of indiscernibles. That is, $d(x,y)=0$ does not necessarily imply that $x$ and $y$ are identical. In the literature, these functions are also termed semi-metrics (e.g., Wells and Williams,, 1975; Burago et al.,, 2001). We denote open and closed balls of radius $r$ by
[…]
respectively. We follow the convention of defining the distance between a point and a set, say $x\in D$ and $A\subset D$, as the infimum distance between them. Formally, $d(x,A)=\inf_{a\in A}d(x,a)$. Finally, the $r$-neighborhood of a set $A\subset D$ is the union of open balls centered at the elements of that set:
$\mathrm{B}_{r}(A)=\bigcup_{a\in A}\mathrm{B}_{r}(a).$
@@
This result serves as a cornerstone for our main findings in Section 3. A proof is provided in Appendix A. For additional background on length spaces, we refer to Appendix B.
@@ 2.1 Conditionally negative definite functions and embeddings
$\sum_{i=1}^{m}\sum_{j=1}^{m}b_{i}b_{j}g(s_{i},s_{j})\leq 0,$ (1)
for $s_{1},\ldots,s_{m}\in D$ and any real numbers $b_{1},\ldots,b_{m}$ such that $\sum_{i}b_{i}=0$ (Rao,, 1984). In contrast, a PD function $r\,:\,D\times D\to\mathbb{R}$ must satisfy
$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}r(s_{i},s_{j})\geq 0,$
for any real numbers $a_{1},\ldots,a_{m}$. Finally, a real function $\varphi$ defined on non-negative real numbers is radial positive definite (RPD) on the pseudometric space $(D,d)$ provided that $\varphi$ is continuous and
$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\varphi(d(s_{i},s_{j}))\geq 0,$ (2)
@@
Denote $\Phi_{p}$ as the class RPD functions on $\mathbb{R}^{p}$. Functions belonging to this class satisfy (Schoenberg,, 1938; Gneiting,, 2013):
@@
Isometric embeddings, a special class of mappings between spaces, are crucial to establish our results. An isometric embedding preserves the distances between points in the original space when mapped to the target space. More formally, a space $(D,d)$ can be isometrically embedded in another space $(D^{\ast},d^{\ast})$ if there exists a map $\phi\,:\,D\to D^{\ast}$ such that
@@
A pseudometric space $(D,d)$ can be isometrically embedded in a Hilbert space if and only if $d^{2}$ is CND.
@@
The proof is presented in Appendix A. Corollary 2.2.1 establishes that the PEXP function with power $\nu\in(0,1]$ is always positive definite on metric spaces equipped with CND metrics (on the original scale).
@@
This result shows that covariance families in $\Phi_{\infty}$, originally defined for Hilbert spaces, remain valid on any domain equipped with a conditionally negative definite pseudometric after a square-root transformation of the distance. This implication allows covariance constructions developed in Euclidean settings to be transferred directly to a wide range of non-Euclidean domains encountered in spatial statistics, including those based on $\alpha$-norms (Curriero,, 2006) and graph-indexed spaces (Anderes et al.,, 2020).
Verifying whether a distance is conditionally negative definite can be challenging in general, but several established results provide useful guidance. For instance, any $\alpha$-norm on $\mathbb{R}^{2}$ is conditionally negative definite (Richards,, 1985). In higher dimensions, that is, $\mathbb{R}^{p}$ with $p\geq 3$, $\alpha$-norms are CND for $\alpha\in(0,2]$. In addition, Theorem 3.6 of (Meckes,, 2013) shows that if a metric $d$ is CND, then so is $d^{\nu}$ for $\nu\in(0,1]$, and that the great-circle distance on round spheres is also CND. In the construction developed in Section 3, however, conditional negative definiteness can be established directly from the structure of the proposed distance, substantially simplifying this verification step for set-indexed domains.
@@ 3 Isotropic positive definite functions for sets
To motivate the proposed distance, we first consider the classical Hausdorff distance, which provides a natural notion of discrepancy between spatial sets but lacks general guarantees for valid covariance construction. For non-empty sets $A_{1},A_{2}\subset D$, the Hausdorff distance is defined as
@@
If no such $r$ exists, set $d_{h}(A_{1},A_{2}):=\infty$. While $d_{h}$ constitutes a valid metric on the class of non-empty, closed, and bounded subsets of $D$ (Arutyunov et al.,, 2016), it presents significant challenges for statistical modeling. Despite its intuitive interpretation in spatial statistics (Godoy et al.,, 2022, 2025), computing the Hausdorff distance is often prohibitively expensive (Knauer et al.,, 2011; Parikh,, 2023). More critically, to the best of our knowledge, constructing valid positive-definite functions based directly on $d_{h}$ remains an open problem. These analytical and computational limitations motivate the ball-Hausdorff distance defined below.
3.1 The ball-Hausdorff distance
@@
We must address the existence and uniqueness of the Chebyshev center to ensure the metric is well-defined. Regarding existence, while the set of centers $\mathcal{E}(A)$ may be empty in arbitrary metric spaces, the Hopf–Rinow–Cohn-Vossen Theorem (Burago et al.,, 2001, Theorem 2.5.28) guarantees that in complete manifolds (such as the sphere and torus), closed and bounded sets are compact. Consequently, the radius function attains its minimum, ensuring $\mathcal{E}(A)$ is non-empty. Regarding uniqueness, the strict convexity of the norm in Euclidean spaces guarantees that $\mathcal{E}(A)$ is a singleton (Amir,, 1978). However, uniqueness is not guaranteed globally in general length spaces. For example, a great circle on $\mathbb{S}^{2}$ admits multiple centers (the poles). Therefore, to ensure validity in non-Euclidean settings, we impose further restrictions on $\mathcal{C}_{D}$. For the unit sphere, this requires a Chebyshev radius less than $\pi/2$ (Bridson and Haefliger,, 1999, Proposition 1.4). This condition satisfied by all spatial configurations in our numerical experiments.
Using the definition 3.1, we define the geometric summary $\mathcal{B}(A)$ as the minimum enclosing ball of $A$, characterized by the center $c(A)$ and radius $R(A)$. Note that, if $A$ is already a ball (or a singleton), it follows trivially that $A=\mathcal{B}(A)$.
@@
The ball–Hausdorff distance inherits basic pseudometric properties from the Hausdorff distance while remaining insensitive to within-set geometry. In particular, for $A_{1},A_{2},A_{3}\in\mathcal{C}_{D}$,
$bh(A_{1},A_{2})$ $\geq 0,$
$bh(A_{1},A_{2})$ $=0,\quad\text{ if }\quad A_{1}=A_{2},$
$bh(A_{1},A_{2})$ $\leq bh(A_{1},A_{3})+bh(A_{2},A_{3}),$
where the first two properties are immediate from the definition and the third follows from the fact that Hausdorff distance is a metric on $\mathcal{C}_{D}$ (Molchanov,, 2005, Appendix C). Note that the distance does not render a metric because $bh(A_{1},A_{2})=0$ does not imply $A_{1}$ and $A_{2}$ are identical. Importantly, the ball-Hausdorff distance reduces to the underlying metric $d$ when applied to singletons. In other words, the proposed distance function between points reduces to $d$, from the original metric space. In Appendix C, we also derive an upper-bound for the ball-Hausdorff distance.
An important question in practice is how to compute the ball-Hausdorff distance. At first sight, it involves computationally expensive operations. For instance, calculating the Hausdorff distance between two balls may not always be trivial (see Section 2 for the problems in calculating the Hausdorff distance). In the special case where the study region lies on the Euclidean plane, a closed-form formula for the Hausdorff distance between circles has been developed (Scitovski and Marošević,, 2015). We extend this formula to length spaces which is general enough to cover most scenarios relevant for spatial statistics applications.
@@
Theorem 3.3 offers both a practical computational method and a clear geometric interpretation. Unlike the standard Hausdorff distance, which depends on extremal point-to-point discrepancies, the ball–Hausdorff distance depends only on relative location and scale through the sets’ Chebyshev centers and radii. In the limiting case where sets have identical radii or are small relative to their separation, the distance reduces to the distance between their centers.
3.2 Isotropic positive definite functions for sets

### v3

# On the Validity of Isotropic Covariance Functions for Set-indexed Random Fields

@@ Abstract
Distances between sets arise naturally when modeling stochastic dependence on collections of spatial supports, including settings with point-referenced and areal observations. However, commonly used constructions of distances on sets, including those derived from the Hausdorff distance, generally fail to be conditionally negative definite, precluding their use in isotropic covariance models. We propose the ball–Hausdorff distance, defined as the Hausdorff distance between the minimum enclosing balls of bounded sets in a metric space. For length spaces, we derive an explicit representation of this distance in terms of the associated centers and radii. We show that the ball–Hausdorff distance is conditionally negative definite whenever the underlying metric is conditionally negative definite. By Schoenberg’s theorem, this implies an isometric embedding into a Hilbert space and guarantees the validity of broad classes of isotropic covariance functions, including the Matérn and powered exponential families, for set-indexed random fields. The construction reduces dependence between sets to low-dimensional geometric summaries, leading to substantial simplifications in covariance evaluation.
Keywords: Change of support, Geostatistics, Hausdorff distance, Length spaces, Spatial data fusion
@@ 1 Introduction
Many modern spatial analyses require defining dependence across observations indexed by spatial units with heterogeneous locations, sizes, and geometries. Several closely related statistical problems arise from this mismatch. Change of support refers to predicting a process indexed on one spatial scale using observations collected on another (Gelfand et al.,, 2001), with downscaling denoting the special case in which information aggregated over coarse spatial units is used to infer values at a finer resolution (Zheng et al.,, 2025). When observations on the same phenomenon are simultaneously available at multiple spatial resolutions, the problem is commonly framed as spatial data fusion (Moraga et al.,, 2017). In addition, response and explanatory variables are often observed on incompatible spatial supports, a setting known as spatial misalignment (Godoy et al., 2026b, ). Although these settings are often treated separately, they expose a common methodological challenge: constructing valid dependence structures on collections of spatial sets rather than on points.
@@
Most existing approaches to multi-resolution spatial data model dependence by treating observations at different supports as aggregations of a common underlying continuous process, which leads to substantial computational and theoretical limitations. Areal measurements are typically represented as integrals of an underlying stochastic process (Fuentes and Raftery,, 2005). However, the implied finite-dimensional probability density functions involve stochastic integrals that are analytically intractable even when the underlying process is a Gaussian Process (GP). Consequently, inferences rely on numerical approximations or Monte Carlo integration (Gelfand et al.,, 2001), which are computationally demanding even for moderately sized datasets and introduce hard-to-quantify bias (Gonçalves and Gamerman,, 2018). An alternative strategy avoids aggregation by defining stochastic processes directly on collections of spatial sets. The Hausdorff distance has been explored as a basis for such constructions (Godoy et al., 2026a, ), but its direct construction does not guarantee positive definite (PD) functions, leaving the existence and validity of the induced process unresolved. This lack of theoretical guarantee, coupled with high computational costs (Knauer et al.,, 2011), necessitates a more robust framework for set-to-set dependence.
@@
Constructing valid PD functions becomes substantially more delicate once the underlying index space departs from Euclidean geometry. While many PD functions are known for $\mathbb{R}^{p}$ equipped with the Euclidean distance (Cressie,, 1993, Ch. 2.5.1), their positive definiteness may not extend to other spaces (Li et al.,, 2023). In fact, a function that is PD on $\mathbb{R}^{1}$ may not remain PD on $\mathbb{R}^{2}$, although the reverse is true (Cressie,, 1993, p. 22). This geometric sensitivity has motivated recent work on characterizing PD covariance functions on non-Euclidean spaces. The particular case of the sphere has been extensively studied (see Porcu et al., (2021) and references therein). Certain covariance functions exhibit the notable property of being PD on both the (p1)(p-1)-dimensional sphere ($\mathbb{S}^{p-1}$) and $p$-dimensional Euclidean space (Porcu et al.,, 2016). More recently, analogous questions have arisen for covariance functions indexed by graphs and their edges (Anderes et al.,, 2020). Despite these advances, there is no general construction of distances on spatial sets that both ensures positive definiteness of isotropic covariance functions and avoids the computational burdens associated with aggregation-based models.
@@
To address this gap, we propose a framework for constructing isotropic covariance functions directly on collections of spatial sets, accommodating point-referenced, areal, and mixed spatial data within a unified formulation. We aim to circumvent the analytical and computational intractability of aggregation-based models while ensuring mathematical validity. To this end, we define a pseudometric, the ball-Hausdorff distance, based on the Hausdorff distance between the minimum enclosing balls of sets. Under mild assumptions, we derive a closed-form expression for this distance, allowing for fast and exact computation. Moreover, the proposed distance is conditionally negative definite (CND) whenever the underlying domain metric is CND. By Schoenberg’s theorem (Schoenberg,, 1938), this implies that the square-root of the ball-Hausdorff distance admits an isometric embedding into a Hilbert space, thereby providing a rich class of valid isotropic covariance functions indexed by spatial sets. As we demonstrate in numerical studies, this theoretical validity is a critical distinction, as the standard Hausdorff distance frequently yields invalid covariance structures in identical settings.
@@ 2 Preliminaries
Let $D$ denote an index set and $d\,:\,D\times D\to[0,\infty)$ a distance function. The pair $(D,d)$ is a pseudometric space if $d$ satisfies all the metric axioms except the identity of indiscernibles. That is, $d(x,y)=0$ does not necessarily imply that $x$ and $y$ are identical. In the literature, these functions are also termed semi-metrics (e.g., Wells and Williams,, 1975; Burago et al.,, 2001). We denote open balls of radius $r$ by $\mathrm{B}_{r}(x)=\{u\in D:d(x,u)<r\}$. We follow the convention of defining the distance between a point and a set, say $x\in D$ and $A\subset D$, as the infimum distance between them. Formally, $d(x,A)=\inf_{a\in A}d(x,a)$. Finally, the $r$-neighborhood of a set $A\subset D$ is the union of open balls centered at the elements of that set: $\mathrm{B}_{r}(A)=\bigcup_{a\in A}\mathrm{B}_{r}(a)$.
@@
Proof.
$d(x,u)\leq d(x,v)+d(v,u)<k+r,$
thus, $u\in\mathrm{B}_{r+k}(x)$.
This result serves as a cornerstone for our main findings in Section 3. For additional background on length spaces, we refer to Appendix A.
@@ 2.1 Conditionally negative definite functions and embeddings
$\sum_{i=1}^{m}\sum_{j=1}^{m}b_{i}b_{j}g(s_{i},s_{j})\leq 0,$
for $s_{1},\ldots,s_{m}\in D$ and any real numbers $b_{1},\ldots,b_{m}$ such that $\sum_{i}b_{i}=0$. In contrast, a real function $\varphi$ defined on non-negative real numbers is radial positive definite (RPD) on the pseudometric space $(D,d)$ provided that $\varphi$ is continuous and
$\sum_{i=1}^{m}\sum_{j=1}^{m}a_{i}a_{j}\varphi(d(s_{i},s_{j}))\geq 0,$
@@
Denote $\Phi_{p}$ as the class of RPD functions on $\mathbb{R}^{p}$. Functions belonging to this class satisfy (Schoenberg,, 1938; Gneiting,, 2013):
@@
Isometric embeddings, a special class of mappings between spaces, are crucial to establish our results. An isometric embedding preserves the distances between points in the original space when mapped to the target space. More formally, a space $(D,d)$ can be isometrically embedded on another space $(D^{\ast},d^{\ast})$ if there exists a map $\phi\,:\,D\to D^{\ast}$ such that
@@
A pseudometric space $(D,d)$ can be isometrically embedded on a Hilbert space if and only if $d^{2}$ is CND.
@@
Proof.
This establishes that $\exp\{-\kappa d^{\nu}\}$ is RPD on $(D,d)$, for $\nu\in(0,1]$. ∎
Corollary 2.2.1 establishes that the PEXP function with power $\nu\in(0,1]$ is always positive definite on metric spaces equipped with CND metrics (on the original scale).
@@
This result shows that covariance families in $\Phi_{\infty}$, originally defined for Hilbert spaces, remain valid on any domain equipped with a conditionally negative definite pseudometric after a square-root transformation of the distance. This implication allows covariance constructions developed in Euclidean settings to be transferred directly to a wide range of non-Euclidean domains encountered in spatial statistics, including those based on $\alpha$-norms (Richards,, 1985), great-circle distance (Meckes,, 2013) and graph-indexed spaces (Anderes et al.,, 2020).
@@ 3 Isotropic positive definite functions for sets
We address the general problem of constructing valid isotropic dependence structures on collections of spatial sets, rather than on point-indexed domains. As a natural starting point, one may consider the classical Hausdorff distance, which provides a natural notion of discrepancy between spatial sets but lacks general guarantees for valid covariance construction. For non-empty sets $A_{1},A_{2}\subset D$, the Hausdorff distance is defined as
@@
If no such $r$ exists, set $d_{h}(A_{1},A_{2}):=\infty$. While $d_{h}$ constitutes a valid metric on the class of non-empty, closed, and bounded subsets of $D$ (Arutyunov et al.,, 2016), it presents significant challenges for statistical modeling. Despite its intuitive interpretation in spatial statistics (Godoy et al.,, 2022; Godoy et al., 2026a, ), computing the Hausdorff distance is often prohibitively expensive (Knauer et al.,, 2011) and, to the best of our knowledge, constructing valid positive-definite functions based directly on $d_{h}$ remains an open problem. These analytical and computational limitations motivate the ball-Hausdorff distance defined below.
@@
We must address the existence and uniqueness of the Chebyshev center to ensure the metric is well-defined. Regarding existence, while the set of centers $\mathcal{E}(A)$ may be empty in arbitrary metric spaces, the Hopf–Rinow–Cohn-Vossen Theorem (Burago et al.,, 2001, Theorem 2.5.28) guarantees that in complete manifolds (such as the sphere and torus), closed and bounded sets are compact. Consequently, the radius function attains its minimum, ensuring $\mathcal{E}(A)$ is non-empty. Regarding uniqueness, the strict convexity of the norm in Euclidean spaces guarantees that $\mathcal{E}(A)$ is a singleton (Amir,, 1978). However, uniqueness is not guaranteed globally in general length spaces. For example, a great circle on $\mathbb{S}^{2}$ admits multiple centers (the poles). Therefore, to ensure validity in non-Euclidean settings, we impose further restrictions on $\mathcal{C}_{D}$. For the unit sphere, this requires a Chebyshev radius less than $\pi/2$ (Bridson and Haefliger,, 1999, Proposition 1.4). This condition is satisfied by all spatial configurations in our numerical experiments.
Using Definition 3.1, we define the geometric summary $\mathcal{B}(A)$ as the minimum enclosing ball of $A$, characterized by the center $c(A)$ and radius $R(A)$. Note that, if $A$ is already a ball (or a singleton), it follows trivially that $A=\mathcal{B}(A)$.
@@
The ball–Hausdorff distance inherits basic pseudometric properties from the Hausdorff distance while remaining insensitive to within-set geometry. In particular, non-negativeness follows immediately from the definition, while the triangle inequality is a consequence of the fact that Hausdorff distance is a metric on $\mathcal{C}_{D}$ (Molchanov,, 2005, Appendix C). Note that the distance does not render a metric because $bh(A_{1},A_{2})=0$ does not imply $A_{1}$ and $A_{2}$ are identical. Importantly, the ball-Hausdorff distance reduces to the underlying metric $d$ when applied to singletons. In other words, the proposed distance function between points reduces to $d$, from the original metric space. In Appendix B, we also derive an upper-bound for the ball-Hausdorff distance.
An important question in practice is how to compute the ball-Hausdorff distance. At first sight, it involves computationally expensive operations. For instance, calculating the Hausdorff distance between two balls may not always be trivial (see Section 2 for the problems in calculating the Hausdorff distance). In the special case where the study region lies on $\mathbb{R}^{2}$, a closed-form formula for the Hausdorff distance between circles has been developed (Scitovski and Marošević,, 2015). We extend this formula to length spaces, which are general enough to cover most scenarios relevant for spatial statistics applications.
@@
Theorem 3.3 offers both a practical computational method and a clear geometric interpretation. Unlike the standard Hausdorff distance, which depends on extremal point-to-point discrepancies, the ball–Hausdorff distance depends only on relative location and scale through the sets’ Chebyshev centers and radii. In the limiting case where sets have identical radii or are small relative to their separation, the distance reduces to the distance between their centers. Despite their structural differences, the ball–Hausdorff and Hausdorff distance may coincide. The corollary below presents one such instance.
Corollary 3.3.1.
In $\mathbb{R}^{1}$, $bh(\cdot,\cdot)\equiv d_{h}(\cdot,\cdot)$.
Proof.
Let $[a,b]$ and $[c,d]$ represent two arbitrary intervals in $\mathbb{R}^{1}$. Then, we have:
$bh([a,b],[c,d])$ $=\frac{1}{2}[\lvert(a+b)-(c+d)\rvert+\lvert(b-a)-(d-c)\rvert$]
$=\frac{1}{2}[\lvert(b-d)+(a-c)\rvert+\lvert(b-d)-(a-c)\rvert$]
[…]
Similarly,
$d_{h}([a,b],[c,d])$ $=\inf\{r\geq 0\,:\,[a,b]\subseteq B_{r}([c,d]),[c,d]\subseteq B_{r}([a,b])\}$
$=\inf\{r\geq 0\,:\,[a,b]\subseteq[c-r,d+r],[c,d]\subseteq[a-r,b+r]\}$
[…]
Therefore,
$bh([a,b],[c,d])=d_{h}([a,b],[c,d]),$
as claimed. ∎